OpenStudy (anonymous):

how do i start to solve an equation when the base isnt the same. 625^3n-2=5^3

OpenStudy (anonymous):

5 ^[4(3n-2)] = 5^3 4(3n-2) = 3 12n -8 =3 12n=11 n = 11/12

OpenStudy (anonymous):

how did u make the bases the same

OpenStudy (anonymous):

bcoz 625 = 5^4